Solve the given matrix using strassen’s matrix multiplication
A matrix is 3 5
5 8
B matrix is 8 9
4 7
1
Expert's answer
2020-04-13T19:39:05-0400

According to Strassen's algorithm, "\begin{pmatrix} a & b \\c & d\end{pmatrix} \cdot \begin{pmatrix} e & f \\g & h\end{pmatrix} = \begin{pmatrix} p_5 + p_4 - p_2 + p_6 & p_1 + p_2 \\p_3 + p_4 & p_1 + p_5 - p_3 - p_7\end{pmatrix}", where


"p_1 = a(f-h)\\
p_2 = (a+b)h\\
p_3 = (c+d)e\\
p_4 = d(g-e)" "p_5 = (a+d)(e+h)\\
p_6 = (b-d)(g+h)\\
p_7 = (a-c)(e+f)"


For our case, "a = 3, b=5, c=5, d=8, e=8, f=9, g=4, h=7", hence calculating, obtain "p_1 = 6, p_2 = 56, p_3 = 104, p_4 = -32, p_5 = 165, p_6 = -33, p_7 = -34".

Substituting these parameters into Strassen's formula, obtain "\begin{pmatrix} 44 & 62 \\ 72 & 101\end{pmatrix}".

The result is easily checked using normal matrix multiplication.